### v1

# bibliography.bib

Abstract
Prior-data fitted networks (PFNs) have recently been proposed as a promising way to train tabular foundation models. PFNs are transformers that are pre-trained on synthetic data generated from a prespecified prior distribution and that enable Bayesian inference through in-context learning. In this paper, we introduce CausalFM, a comprehensive framework for training PFN-based foundation models in various causal inference settings. First, we formalize the construction of Bayesian priors for causal inference based on structural causal models (SCMs) in a principled way and derive necessary criteria for the validity of such priors. Building on this, we propose a novel family of prior distributions using causality-inspired Bayesian neural networks that enable CausalFM to perform Bayesian causal inference in various settings, including back-door, front-door, and instrumental variable adjustment. Finally, we instantiate CausalFM and explicitly train a foundation model for estimating conditional average treatment effects (CATEs) using back-door adjustment. We show that CausalFM performs competitively for CATE estimation using various synthetic and semi-synthetic benchmarks. In sum, our framework can be used as a general recipe to train foundation models for various causal inference settings. In contrast to the current state-of-the-art in causal inference, CausalFM offers a novel paradigm with the potential to fundamentally change how practitioners perform causal inference in medicine, economics, and other disciplines.
1 Introduction
Causal inference is a cornerstone of empirical research in disciplines such as economics [Angrist.1990, Imbens.1994], medicine [Feuerriegel.2024, Glass.2013], and marketing [Varian.2016]. It enables the estimation of causal effects from observational and randomized data, which is essential for reliable decision-making. In personalized medicine, for instance, it supports identifying the most effective treatment by predicting patient outcomes under different therapeutic options.
In recent years, machine learning, and especially deep learning methods, have gained significant traction in causal inference (e.g., [Curth.2021, Melnychuk.2022, Schweisthal.2023, Shalit.2017, Shi.2019]). These methods offer several advantages for causal effect estimation in practice, including the ability to handle large-scale and high-dimensional datasets with complex confounding structures and to model heterogeneity of causal effects [Feuerriegel.2024]. However, most existing approaches require retraining a model for each new causal task or dataset. To this end, existing approaches lack the flexibility to perform test-time inference without additional retraining, which limits their practicality in real-world settings.
Meanwhile, foundation models have emerged as a transformative paradigm in machine learning [devlin2018bert, touvron2023llama, touvron2023llama2, lahat2024assessing], which offer a key advantage in that they allow for flexible, test-time inference without retraining. These models are pre-trained on large datasets and can generalize across tasks and domains. Examples include large language models (LLMs) in natural language processing and vision transformers in computer vision. However, this paradigm shift toward test-time inference has not yet had a comparable impact on causal inference. Most current approaches in causal machine learning still rely on specialized models tailored to specific tasks, requiring practitioners to manually select, train, and validate appropriate estimation methods for each new dataset.
In this paper, we propose a change to the paradigm for causal inference based on the idea of foundation models trained for tabular causal inference. For this, we build on the recently proposed prior-data fitted networks (PFNs) [Muller.2022, Hollmann.2023], which are transformers pre-trained on purely synthetic datasets generated from a prespecified prior distribution. PFNs enable Bayesian inference purely through in-context learning, allowing for flexible and efficient predictions without requiring additional training at inference time [Nagler.2023]. While recent works have demonstrated the effectiveness of tabular foundation models based on PFNs for various tasks, only two concurrent works have proposed PFNs tailored for causal inference [Balazadeh.2025, Robertson.2025]. However, these are either restricted to specific causal inference settings (namely, only back-door adjustment) or do not offer identifiability guarantees.
We introduce CausalFM, a comprehensive framework for training PFN-based foundation models for various causal inference settings. For this purpose, we introduce the CausalFM prior: a novel family of prior distributions based on structural causal models respecting the underlying causal inference problem at hand. We first formalize and derive necessary criteria on how to construct such SCM-based priors for causal inference in principle. Then, we propose a concrete instantiation using Bayesian neural networks and provide a learning algorithm that leverages the SCM’s ability to simulate counterfactual data and perform consistent Bayesian inference in a wide range of causal inference settings.
Compared to classical causal inference methods, foundation models trained based on our CausalFM offer the following advantages: (i) There is no need for additional training at inference time as our CausalFM performs inference entirely through in-context learning, enabling fast and flexible deployment across new datasets. (ii) The Bayesian nature of our CausalFM provides principled uncertainty quantification, which is critical for downstream decision-making and for detecting situations with poor treatment overlap. (iii) The model automatically learns to select among alternative adjustment strategies such as back-door, front-door, or instrumental variable (IV) adjustment based on the data distribution and task at hand. For instance, in the presence of a weak instrument, the model may automatically choose to avoid IV adjustment in favor of a more robust alternative, thereby improving estimation performance and reducing variance. (iv) Our CausalFM builds upon rigorous identifiability guarantees to ensure asympotically valid inferences.
Our contributions [^1] Code is available at https://github.com/yccm/CausalFM. are: (1) We formalize the constructions of priors based on structural causal models (SCMs) for Bayesian causal inference and derive necessary conditions for their validity. (2) We propose an explicit CausalFM prior based on Bayesian neural networks that are compatible with the structure of the causal inference problem at hand. We also propose a learning algorithm to train PFNs for causal inference problems that leverages our CausalFM prior to simulate counterfactuals to mitigate the fundamental problem of causal inference. (3) We propose a concrete instantiation of our framework: we train a PFN for estimating conditional average treatment effects (CATEs) via back-door adjustment. We show empirically that CausalFM performs competitively and outperforms current state-of-the-art CATE estimators on a variety of benchmarks.
2 Related Work
Prior-data-fitted networks (PFNs) as tabular foundation models.
Foundation models are an emerging paradigm that has revolutionized machine learning for various data modalities, particularly language and vision tasks [devlin2018bert, touvron2023llama, touvron2023llama2, lahat2024assessing]. The same paradigm is now being explored for tabular data—a modality that underpins the large majority of analyses in science and business [vanBreugel.2024].
A powerful approach is given by prior-data-fitted networks (PFNs) [Muller.2022], which are large transformers trained exclusively on synthetic data sampled from a user-specified prior distribution. Then, at test time, it performs Bayesian inference through in-context learning by approximating the posterior predictive distribution. TabPFN [Hollmann.2023, Hollmann.2025] scaled this idea by pairing the transformer with a Bayesian neural-network prior over structural causal models (SCMs) and demonstrating state-of-the-art performance on various tabular benchmarks. Subsequent work extended PFNs to time–series forecasting [Hoo.2025] and analyzed their in-context learning abilities theoretically [Nagler.2023]. Critically, all existing PFNs are trained only for predictive tasks and do not target causal estimands; they therefore are not designed for causal inference of treatment effects, which is the goal of our paper.
Causal inference with machine learning.
Causal inference, such as the estimation of average treatment effects, originates from fields like econometrics [Imbens.1994, Angrist.1990], statistics [vanderLaan.2006], and epidemiology [Robins.1986, Robins.1994]. More recently, machine learning has focused on estimating heterogeneous effects to support personalized decision-making.
One line of work are frequentist methods, which build on semiparametric theory and efficient influence functions [Robins.1994b, Robins.1999], yielding model-agnostic estimators that are doubly robust and Neyman-orthogonal [vanderLaan.2006b, Chernozhukov.2018, Nie.2021, Foster.2023, Kennedy.2023b]. Another line of work builds upon specific machine learning methods/ architectures such as regression trees [Wager.2018] or neural networks [Johansson.2016, Shalit.2017, Shi.2019] and adopts them to causal inference. Bayesian alternatives include Bayesian additive regression trees [Hahn.2020] or Gaussian-process counterfactual regression [Alaa.2017]. However, all of the existing estimators above must be retrained for every new dataset and causal setting. In contrast, our CausalFM allows to pre-train models to approximate Bayesian causal inference via in-context learning.
Amortized causal inference.
Several recent papers pre-train large neural networks on synthetic data so that they can solve causal tasks via in-context learning during inference time. Examples include causal discovery [Mahajan.2025], ATE estimation under unconfoundedness [Zhang.2024b], zero-shot- and few-shot learning [Nilforoshan.2023, Iwata.2023], and reinforcement-learning [Lee.2023]. These methods validate the feasibility of treating causal inference as an in-context learning problem but remain restricted to specific causal inference settings – which typically do not allow to accommodate unobserved confounding.
Black-box causal inference (BBCI) [Bynum.2025] takes a step towards pertaining models on synthetic data to perform causal inference in a variety of settings. However, their approach is different as compared to our CausalFM as follows: (i) BBCI does not build upon a Bayesian framework. In contrast, building upon PFNs allows us to perform approximate Bayesian causal inference and thus provide rigorous uncertainty quantification for treatment effects. (ii) The proposed data-generating processes in BBCI are not tailored for high-dimensional causal inference settings (as the authors mention in their Sec. 7). In contrast, our CausalFM prior leverages Bayesian neural networks inspired by TabPFN [Hollmann.2023] to create high-dimensional SCM-based prior distributions. (iii) Beyond proposing a new method, we provide novel formalizations and theoretical results of constructing valid SCM-based priors for Bayesian causal inference.
Foundation models for causal inference: We are aware of only two works published concurrently to this paper that propose PFN-based models for causal inference, but each with clear limitations: (i) [Balazadeh.2025] proposes a PFN similar to ours, but it is restricted to only back-door adjustment, i.e., imposes the unconfoundedness assumption throughout their paper. In contrast, we propose a framework for constructing PFN-based foundation models for a large class of causal inference problems, including both front-door adjustment and instrumental variable settings with unobserved confounding. As a consequence, our Bayesian prior construction for training the PFN is different and allows to sample various causal inference settings. (ii) [Robertson.2025] proposes to train a single PFN on various different causal inference settings without providing identifiability assumptions to the model. We will show later that the approach of [Robertson.2025] has a crucial drawback: because the causal quantity of interest is not identified, the PFN learns a posterior that may never concentrate around the true causal quantity, thus leading to asymptotically biased estimators. In contrast, we propose to train separate PFNs for different identifiable causal inference settings. As such, we follow established philosophy in causal inference that separates identifiability and estimation steps [Pearl.2009]: the identifiability step should be established by the practitioner using domain knowledge (e.g., establishing whether a certain variable is a valid instrument), while the estimation step can be treated as a purely statistical learning problem and thus can be handed over to a pre-trained foundation model.
3 Problem setup
3.1 Background on PFNs
In tabular prediction problems, one considers a population $(X,Y)\sim\mathbb{P}\in\mathcal{P}$. Given a finite sample $\mathcal{D}_{n}\sim\mathbb{P}$ of size $n$, the goal is to estimate the conditional distribution $\mathbb{P}(Y=y\mid X=x)$. PFNs formulate this task in a Bayesian non-parametric way by placing a prior distribution $\Pi$ on $\mathcal{P}$, i.e., a prior over data-generating distributions [Muller.2022, Nagler.2023]. Sampling proceeds hierarchically via
$\mathbb{P}\sim\Pi,\qquad(X_{i},Y_{i})\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathbb{P}.$ (1)
Then, Bayes’ rule yields the posterior distribution
$\Pi(\mathbb{P}\mid\mathcal{D}_{n})\;\propto\;\Pi(\mathcal{D}_{n}\mid\mathbb{P})\,\Pi(\mathbb{P}),$ (2)
where $\Pi(\mathcal{D}_{n}\mid\mathbb{P})$ is the likelihood of the sample $\mathcal{D}_{n}$ under $\mathbb{P}$ and $\propto$ denotes proportionality up to a multiplicative constant. The corresponding posterior-predictive distribution is defined as
$\Pi(Y\mid\mathcal{D}_{n},x)=\int\mathbb{P}(Y\mid X=x)\,\Pi(\mathbb{P}\mid\mathcal{D}_{n})\,\mathop{}\!\mathrm{d}\mathbb{P},$ (3)
which is the probability of $Y$ given test point $x$ and observed data $\mathcal{D}_{n}$.
PFNs are neural networks $q_{\theta}(Y\mid\mathcal{D}_{n},x)$ that parameterize the family of predictive posterior distributions with trainable parameters $\theta$. That is, PFNs map the entire dataset $\mathcal{D}_{n}$ and a query $x$ to a distribution over $\mathcal{Y}$. In terms of architecture, PFNs are permutation-equivariant transformers [Vaswani.2017] as they allow for scalable training and leverage the attention mechanism to effectively extract information from $\mathcal{D}_{n}$.
PFNs are trained by minimizing the negative log-likelihood loss
$\mathcal{L}(\theta)=\mathbb{E}_{N\sim\Pi_{N}}\!\bigl{[}\mathbb{E}_{\mathbb{P}\sim\Pi}\bigl{[}-\log q_{\theta}(Y\mid X,\mathcal{D}_{N})\bigr{]}\bigr{]},$ (4)
where $\Pi_{N}$ is a prior on the sample sizes. In practice, we sample a sample size $N_{j}\sim\Pi_{N}$, a probability distribution $\mathbb{P}_{j}\sim\Pi$, a dataset $\mathcal{D}_{N_{j}}^{j}\sim\mathbb{P}_{j}$, and test points $(x_{j},y_{j})\sim\mathbb{P}_{j}$ and then approximate the PFN loss via
^$\hat{\mathcal{L}}(\theta)=\sum_{j}\bigl{[}-\log q_{\theta}(y_{j}\mid\mathcal{D}_{N_{j}}^{j},x_{j})\bigr{]},$ (5)
which is consistent for the exact posterior-predictive under regularity conditions [Nagler.2023]. Note that all training data are synthetic, i.e., sampled from the prior $\Pi$. Furthermore, the trained PFN can be deployed on arbitrary real datasets without further training.
3.2 Task
In this paper, we aim to extend PFNs to causal inference. Here, the main challenge is that the object of interest is an interventional [^2] Causal literature often distinguishes between interventional and counterfactual distributions. This is not relevant for the methods of our paper, and we thus use interventional distribution as an umbrella term that incorporates both. distribution $\mathbb{P}_{\mathrm{int}}$, yet we only observe data $\mathcal{D}_{n}\sim\mathbb{P}_{\mathrm{obs}}$ from a potentially different observational distribution [Pearl.2009].
Motivation.
As an illustrative example, we consider a standard causal inference setting where the data comprise $(X,A,Y)\sim\mathbb{P}_{\mathrm{obs}}$, where $X$ are patient covariates, $A$ is a treatment and $Y$ an outcome of interest [vanderLaan.2006, Shalit.2017]. For example, in medicine, $X$ may contain treatment history or demographic attributes $A$ may be a medical treatment, and $Y$ a health outcome. Following the potential outcome framework [Rubin.1974], let $Y(a)$ denote the outcome that would realize under the intervention $\mathrm{do}(A=a)$. The interventional distribution is thus over $(X,A,Y(1)-Y(0))\sim\mathbb{P}_{\mathrm{int}}$, and a common target functional is the conditional average treatment effect (CATE) $Q(x)\;=\;\mathbb{E}\bigl{[}Y(1)-Y(0)\mid X=x\bigr{]}$ [Wager.2018]. The CATE quantifies the expected benefit of providing treatment given the patient’s covariates.
Identifiability. To estimate CATE from observational data, we need to impose identifiability assumptions, which link the observational and the interventional distributions and allow us to express $Q$ as a functional of the observational distribution [Rosenbaum.1983]. These are (i) consistency: $Y(A)=Y$, (ii) positivity: $\mathbb{P}_{\mathrm{obs}}(A=1\mid X=x)>0$, and (iii) Unconfoundedness: $Y(1),Y(0)\perp\!\!\!\perp A\mid X$ in $\mathbb{P}_{\mathrm{int}}$.
Instead of Assumption (iii), weaker assumptions are often made (e.g., assuming an instrumental variable while allowing for unconfoundedness). We unify these different assumptions and causal queries in the following.
Generalized causal inference setting. In the following, we provide a generalized definition of a causal inference setting, that allows us to reason about arbitrary causal inference settings and provide generalized statements beyond the standard example above.
Definition 3.1.
We define a causal inference setting as a tuple
$\mathcal{C}\;=\;\bigl{(}Z,\mathcal{P}_{\mathrm{obs}}\times\mathcal{P}_{\mathrm{int}},\,Q\bigr{)},$ (6)
where $Z$ collects the observed variables (and contains at least $A$ and $Y$); $(\mathbb{P}_{\mathrm{obs}},\mathbb{P}_{\mathrm{int}})\in\mathcal{P}_{\mathrm{obs}}\times\mathcal{P}_{\mathrm{int}}$ are paired observational/interventional distributions over $Z$ that correspond to an intervention on $A$; and $Q(\mathbb{P}_{\mathrm{int}})$ is a causal query that is identifiable, i.e. there exists a measurable functional $\bar{Q}$ such that
$Q(\mathbb{P}_{\mathrm{int}})\;=\;\bar{Q}(\mathbb{P}_{\mathrm{obs}})\quad\text{for all }(\mathbb{P}_{\mathrm{obs}},\mathbb{P}_{\mathrm{int}})\in\mathcal{P}_{\mathrm{obs}}\times\mathcal{P}_{\mathrm{int}}.$ (7)
3.2.1 Running examples
[…]
$\bar{Q}(\mathbb{P}_{\mathrm{obs}})(x)\;=\;\mathbb{E}_{\mathbb{P}_{\mathrm{obs}}}[Y\mid A=1,X=x]\;-\;\mathbb{E}_{\mathbb{P}_{\mathrm{obs}}}[Y\mid A=0,X=x].$ (8)
[…]
Identifiability assumptions. We restrict to pairs $(\mathbb{P}_{\mathrm{obs}},\mathbb{P}_{\mathrm{int}})$ that satisfy (i) consistency: $Y=Y(A,M)$ and $M=M(A)$; (ii) positivity: $\mathbb{P}_{\mathrm{obs}}(A=a\mid X=x)>0$ and $\mathbb{P}_{\mathrm{obs}}(M=m\mid A=a,X=x)>0$; and (iii) front-door criterion $M(a)\;\perp\!\!\!\perp\;A\mid X=x,\quad\text{and}\quad Y(a^{\prime},m)\;\perp\!\!\!\perp\;M\mid A=a^{\prime},\,X=x$. Under these assumptions, the CATE is identified via
$\widetilde{Q}(\mathbb{P}_{\mathrm{obs}})(x)$ $\;=\;\sum_{m\in\mathcal{M}}\mathbb{E}_{\mathbb{P}_{\mathrm{obs}}}\!\bigl{[}Y\mid A=1,M=m,X=x\bigr{]}\,\mathbb{P}_{\mathrm{obs}}(A=1\mid X=x)\,\mathbb{P}_{\mathrm{obs}}(M=m\mid A=1,X=x)$ (9)
$-\sum_{m\in\mathcal{M}}\mathbb{E}_{\mathbb{P}_{\mathrm{obs}}}\!\bigl{[}Y\mid A=0,M=m,X=x\bigr{]}\,\mathbb{P}_{\mathrm{obs}}(A=0\mid X=x)\,\mathbb{P}_{\mathrm{obs}}(M=m\mid A=0,X=x),$ (10)
which is the conditional version of Pearl’s front-door adjustment [Pearl.2009], thereby proving that $Q(\mathbb{P}_{\mathrm{int}})=\widetilde{Q}(\mathbb{P}_{\mathrm{obs}})$ for all admissible pairs $(\mathbb{P}_{\mathrm{obs}},\mathbb{P}_{\mathrm{int}})$.
[…]
$Q(\mathbb{P}_{\mathrm{int}})(x,a)\;=\;\mathbb{E}\!\bigl{[}Y(a)\mid X=x\bigr{]},$ (11)
Identifiability assumptions. We restrict to pairs $(\mathbb{P}_{\mathrm{obs}},\mathbb{P}_{\mathrm{int}})$ that satisfy the following conditions [Newey.2003]: (i): Additive structural equation: $Y=f(X,A)+g(X,U)$, with an unknown functions $f$ and $g$ and unobserved confounder $U$, implying that $Y$ does not directly depend on $Z$; (ii) Independence: $U\perp\!\!\!\perp Z\mid X$; (iii) Relevance: $\mathbb{P}_{\mathrm{obs}}(A\mid X=x,Z=z)>0$ is non-constant in $z$; and (iv) Completeness: For every measurable $g$, if $\mathbb{E}_{\mathbb{P}_{\mathrm{obs}}}[g(A,x)\mid X=x,Z=z]=0$ for all $z$, then $g(A,x)=0$ almost surely in $\mathbb{P}_{\mathrm{obs}}$. Then, the CAPO can be written as $Q(\mathbb{P}_{\mathrm{int}})(x,a)=f(x,a)+\mathbb{E}_{\mathbb{P}_{\mathrm{int}}}[g(x,U)\mid X=x]$ and is identified via the integral equation
$\mathbb{E}_{\mathbb{P}_{\mathrm{obs}}}[Y\mid X=x,Z=z]$ $=\mathbb{E}_{\mathbb{P}_{\mathrm{obs}}}[f(X,A)\mid X=x,Z=z]+\mathbb{E}_{\mathbb{P}_{\mathrm{int}}}[g(X,U)\mid X=x]$ (12)
$=\mathbb{E}_{\mathbb{P}_{\mathrm{obs}}}[Q(\mathbb{P}_{\mathrm{int}})(X,A)\mid X=x,Z=z],$ (13)
which has a unique solution $\bar{Q}(\mathbb{P}_{\mathrm{obs}})(x,a)$ due to the completeness condition (iv) [Newey.2003, Hartford.2017].
Research question: PFNs have shown to be an effective way to construct tabular foundation models. However, a causal inference setting $\mathcal{C}$ comes with additional challenges, such as the distinction between observational and interventional distribution ans well as identifiability assumptions. Our main research question is:
Research question How can we train PFNs for a causal inference setting $\mathcal{C}$ that provides a Bayesian estimator of $Q(\mathbb{P}_{\mathrm{int}})$ given an observational dataset $\mathcal{D}_{n}\sim\mathbb{P}_{\mathrm{obs}}$ and some context (e.g., values $x$ or $a$)?
Below, we introduce our CausalML consisting of (1) appropriate priors (see Section 4) and (2) a training algorithm for the underlying PNFS (see Section 5).
4 CausalFM: Priors
The first step to derive our CausalML (and thus train a PFN) is to construct a valid prior distribution $\Pi$ that allows for Bayesian inference and approximating the predictive posterior distribution as in Eq. (5). In this section, we discuss how to construct such priors for a causal inference setting.
4.1 Motivation for SCM-based priors
We motivate our approach to constructing priors with a normally distributed IV setting. That is, we consider Example 3 from above (ignoring $X$) where $(Z,A,Y)\sim\mathbb{P}_{\mathrm{obs}}$ is a multivariate normal distribution with mean zero. A straightforward way would be to put a prior on the parameters (covariance matrix) of this distribution, thus inducing a prior $\Pi$ on the observational distribution $\mathbb{P}_{\mathrm{obs}}$. If the posterior $\Pi(\mathbb{P}_{\mathrm{obs}}\mid\mathcal{D}_{n})\to\mathbb{P}_{\mathrm{obs}}^{\ast}$ converges to the ground-truth observational distribution $\mathbb{P}_{\mathrm{obs}}^{\ast}$ (i.e., satisfying a Bernstein-von-Mises theorem), we can obtain a consistent Bayesian estimator of our causal query via $\bar{Q}(\Pi(\mathbb{P}_{\mathrm{obs}}\mid\mathcal{D}_{n}))$. Accordingly, we could train a PFN $q_{\theta}(Y\mid\mathcal{D}_{n},a)$ with the loss in Eq. (5) and estimate the CAPO via $\bar{Q}(q_{\theta}(Y\mid\mathcal{D}_{n},a))$.
However, the above approach has drawbacks: (i) It requires knowledge of the identification formula $\bar{Q}$, which must be determined in a case-by-case basis depending on the causal inference setting $\mathcal{C}$ at hand. This can be tidious or even hard to compute in practice (e.g., solving the integral equation Eq. (12)). (ii) Putting a prior on $\mathbb{P}_{\mathrm{obs}}$ makes it harder control the distribution of the causal query $Q$ directly. It has been shown in the literature that this can lead to prior misspecification for Bayesian causal inference or slowly converging posterior distributions.
As a remedy, we propose to construct prior distributions over structural causal models (SCMs) as done in, e.g., [Hollmann.2023]. As a toy example, let us consider the following class of SCMs:
$U\sim\mathcal{N}(0,1),\,\epsilon_{Z}\sim\mathcal{N}(0,1),\,\epsilon_{A}\sim\mathcal{N}(0,1),\,\epsilon_{Y}\sim\mathcal{N}(0,1),$ (14)
$Z=\alpha\epsilon_{Z},+\kappa U\,A=\beta Z+\delta\epsilon_{A}+\gamma U,\,Y=\zeta A+\eta U+\theta\epsilon_{Y},$ (15)
where $U$ is an unobserved confounder between $A$ and $Y$, and $\epsilon_{Z}$, $\epsilon_{A}$, $\epsilon_{Y}$ are noise variables, and $\alpha$, $\beta$, $\gamma$, $\delta$, $\zeta$, $\eta$, and $\theta$ are scalars describing the functional dependences between observed and noise variables. Our causal query is $Q(\mathbb{P}_{\mathrm{int}})=\mathbb{E}[Y(1)]=\zeta$.
General approach. The class of SCMs above is compatible with the linear IV setting whenever it holds that $\kappa=0$ (independence assumption from Example 3). Hence, we can specify a prior distribution over this class of SCMs by specifying a distribution $\Pi$ over $(\alpha,\beta,\gamma,\delta,\zeta,\eta,\theta)$ and setting $\kappa=0$. Note that this automatically specifies a distribution over $\mathbb{P}_{\mathrm{obs}}$ (by sampling from the SCM) and $\mathbb{P}_{\mathrm{int}}$ (by intervening and setting $A=1$ in the SCM). Interestingly, this addresses the two drawbacks of observational priors from above as follows: (i) During the PFN training we can sample $\mathcal{D}_{n}\sim\mathbb{P}_{\mathrm{obs}}$ and $y(1)\sim\mathbb{P}_{\mathrm{int}}$ and thus fit $q_{\theta}(y(1)\mid\mathcal{D}_{n})$ for the interventional outcome (see Sec. 5 for details). For estimating the causal query we can thus use $Q(q_{\theta}(y(1)\mid\mathcal{D}_{n}))$ and do not need access to the potentially unknown $\bar{Q}$. (ii) We can directly control the marginal prior distribution of $\zeta$, thus remedying the above drawbacks and allowing us more control to incorporate prior information of our causal query.
Adding identifiability assumptions to the prior. A key question is whether we should actually impose the identifiability assumption $\kappa=0$ when constructing a prior. A different approach would be to also put a prior on $\kappa$, thus taking account the possibility of identifiability violations in the prior. Such an approach has been proposed by [Robertson.2025], where the authors construct a prior over many possible causal inference settings simultaneously. However, as we show in the following, this would make consistent Bayesian estimation of the causal query of interest impossible.
Lemma 4.1.
Let $S^{\ast}=(\alpha^{\ast},\beta^{\ast},\delta^{\ast},\gamma^{\ast},\zeta^{\ast},\eta^{\ast},\theta^{\ast},\kappa^{\ast}=0)$ be an identified ground-truth SCM. Then for any causal target $\zeta\neq\zeta^{\ast}$ there exists another SCM $S=(\alpha,\beta,\gamma,\delta,\zeta,\eta,\theta,\kappa)$ with $\kappa\neq 0$ that induces the same observational distribution as $S^{\ast}$.
Proof.
See Appendix A. ∎
Lemma 4.1 has an important consequence: if our prior $\Pi$ puts positive probability mass on all possible combinations of $(\alpha,\beta,\gamma,\delta,\zeta,\eta,\theta,\kappa)$, the corresponding posterior $\Pi(\cdot\mid\mathcal{D}_{n})$ will even for $n\to\infty$ put positive probability mass on any $\zeta\in\mathbb{R}$, thus being completely non-informative about the causal target quantity. As a consequence, any Bayesian point estimator using such a prior (e.g., as in under the approach [Robertson.2025]) will be asymptotically biased.
In contrast, we present a different approach to circumvent the above problems: namely, we propose to construct PFN-priors that incorporate assumptions that allow for identifiability of the causal target quantity (e.g., setting $\kappa=0$ in the above example). As such, we follow established philosophy in causal inference that separates identifiability and estimation steps [Pearl.2009]: the identifiability step should be established by the practitioner using domain knowledge (e.g., establishing whether a certain variable is a valid instrument). Once identifiability has been established, we can use Bayesian modeling and PFN-based models for the estimation step.
Which noise variables to model? A key question that remains is what classes of SCMs we can use to specify priors for the causal inference setting $\mathcal{C}$ at hand. Indeed, the class of SCMs is non-unique: it turns out that it is not necessary to specify both noise variables $\epsilon_{A}$ and $\epsilon_{Y}$ as the next lemma shows.
Lemma 4.2.
Let $\mathcal{S}^{\ast}=(\alpha^{\ast},\beta^{\ast},\gamma^{\ast},\delta^{\ast},\zeta^{\ast},\eta^{\ast},\theta^{\ast})$ be a fixed SCM from the above class with $\mathrm{Var}^{\ast}(A\mid z)>0$ and $\mathrm{Var}^{\ast}(Y\mid a)>0$ for all $z,a$. Then, there exist unique SCMs $\mathcal{S}_{1}=(\alpha_{1},\beta_{1},\gamma_{1},\delta_{1}=0,\zeta_{1},\eta_{1},\theta_{1})$ and $\mathcal{S}_{2}=(\alpha_{2},\beta_{2},\gamma_{2},\delta_{2},\zeta_{2},\eta_{2},\theta_{2}=0)$ that induce the same observational distribution as $\mathcal{S}^{\ast}$ and thus the same causal query $\zeta_{1}=\zeta_{2}=\zeta^{\ast}$. However, whenever it holds that both $\delta=0$ and $\theta=0$, there exists an SCM $\mathcal{S}^{\ast}$ for which $\zeta\neq\zeta^{\ast}$.
Proof.
See Appendix A. ∎
Lemma 4.2 implies that it suffices to specify priors over SCMs without either treatment noise $\epsilon_{A}$ or outcome noise $\epsilon_{Y}$. However, if we remove both, there exist interventional distributions for which the prior will never put probability mass on the ground-truth causal query, rendering Bayesian inference inconsistent. In the following, we generalize this result to arbitrary SCMs and causal inference settings.
4.2 Formalizing SCM-based priors
In the previous example, we argued why sampling entire structural causal models (SCMs) can yield priors that are better aligned with causal queries than priors defined directly on observational distributions. We now provide a general formalism for such priors.
Definition 4.3 (SCMs [Pearl.2009]).
[…]
Every SCM induces a unique directed acyclic graph (DAG), $\mathcal{G}^{\mathcal{S}}$ by defining mapping of the parents $pi(Z_{i})$ to $Z_{i}$ with directed edges. We distinguish two types of latent variables $U_{i}$ in $\mathcal{G}^{\mathcal{S}}$: $U_{i}$ is an unobserved confounder if it is the parent of both $A$ and $Y$, otherwise, we call it a noise variable.
Intuitively, an SCM is a simulator: we can draw latent variables $U\sim\mathbb{P}$ and pass them through structural functions $f$, resulting in an induced observational distribution $\mathbb{P}_{\mathrm{obs}}^{\mathcal{S}}$ over $Z$. At the same time, we can modify the SCM by intevening on a variable via $do(A=a)$, i.e., fixing the variable and then sampling from the SCM mechanism. This induces a corresponding interventional distribution $\mathbb{P}_{\mathrm{int}}^{\mathcal{S}}$. We call an SCM $\mathcal{S}$ compatible with a causal inference setting $\mathcal{C}$, if
$(\mathbb{P}_{\mathrm{obs}}^{\mathcal{S}},\mathbb{P}_{\mathrm{int}}^{\mathcal{S}})\in\mathcal{P}_{\mathrm{obs}}\times\mathcal{P}_{\mathrm{int}}.$ (16)
Priors over SCMs.
We now define prior distribution over classes of SCMs that enable Bayesian causal inference.
Definition 4.4 ($\mathcal{C}$-SCM-Priors).
A $\mathcal{C}$-SCM-Prior is any probability measure $\Pi(\mathcal{S})$ that puts all its mass on SCMs compatible with $\mathcal{C}$. Via the map $\mathcal{S}\mapsto(\mathbb{P}_{\mathrm{obs}}^{\mathcal{S}},\mathbb{P}_{\mathrm{int}}^{\mathcal{S}})$ every such prior induces a distribution $\Pi\bigl{(}(\mathbb{P}_{\mathrm{obs}},\mathbb{P}_{\mathrm{int}})\bigr{)}$ on $\mathcal{P}_{\mathrm{obs}}\times\mathcal{P}_{\mathrm{int}}$.
Sampling from $\Pi$ therefore amounts to sampling a random latent distribution $\mathbb{P}$ over $U$ as well as random functional assignments $f$. These can them ne used to intern sample an observational dataset $\mathcal{D}_{n}$, i.e., there is a well defined likelihood $\Pi(\mathcal{D}_{n}\mid\mathcal{S})$ induced by $\mathcal{S}$. As a consequence, we can define the posterior distribution over SCMs via
$\Pi(\mathcal{S}\mid\mathcal{D}_{n})\propto\Pi(\mathcal{D}_{n}\mid\mathcal{S})\Pi(\mathcal{S}),$ (17)
where $\propto$ denotes proportionality up to a normalization constant.
Well-specified priors.
The question is now how we should design our prior $\Pi$ such that the induced posterior $\Pi(\mathcal{S}\mid\mathcal{D}_{n})$ allows for valid Bayesian causal inference. We now define a key desirable property of such priors.
Definition 4.5 (Well-specified priors).
A prior $\Pi(\mathcal{S})$ is well-specified for $\mathcal{C}$ if, for every true pair $(\mathbb{P}_{\mathrm{obs}}^{\ast},\mathbb{P}_{\mathrm{int}}^{\ast})$ and every sample $\mathcal{D}_{n}\!\sim\!\mathbb{P}_{\mathrm{obs}}^{\ast}$,
$Q\!\Bigl{(}\int\mathbb{P}_{\mathrm{int}}^{\mathcal{S}}\,\Pi(\mathcal{S}\mid\mathcal{D}_{n})\,\mathop{}\!\mathrm{d}\mathcal{S}\Bigr{)}\;\longrightarrow\;Q(\mathbb{P}_{\mathrm{int}}^{\ast}),\quad n\to\infty.$ (18)
In other words, a well-specified prior ensures that the causal query $Q$ evaluated on the posterior-predictive interventional distribution (PPID) $\int\mathbb{P}_{\mathrm{int}}^{\mathcal{S}}\,\Pi(\mathcal{S}\mid\mathcal{D}_{n})\,\mathop{}\!\mathrm{d}\mathcal{S}$ is a consistent estimator of the causal target. If we were able to train a PFN to approximate the PPID of a well-specified prior, we are sure that we can apply $Q$ on this distribution and obtain a consistent estimator.
Cluster-DAGs.
Because an SCM-prior induces a distribution over possibly many DAGs, we compress them into a shared structure.
Definition 4.6 (C-DAGs [Anand.2023]).
Given variables $\left(Z,U\right)$, a Cluster-DAG (C-DAG) is a DAG on clusters $C_{1},\dots,C_{k}$ which are disjoint subsets of $(Z,U)$.
Each $\mathcal{C}$-SCM-Prior induces a unique C-DAG via the following algorithm: (i) draw an edge $C_{i}\!\to\!C_{j}$ whenever any SCMs $\mathcal{S}$ with $\Pi(\mathcal{S})>0$ contains some arrow from any node of $C_{i}$ to any node of $C_{j}$ and no SCM $\mathcal{S}$ with $\Pi(\mathcal{S})>0$ contains some arrow from any node of $C_{j}$ to any node of $C_{i}$; (ii) merge $C_{i}$ and $C_{j}$ whenever both directions occur across SCMs $\mathcal{S}$ with $\Pi(\mathcal{S})>0$.
Every induced C-DAG inherits the well-specified proprty from its prior.
Definition 4.7 (Well-specified C-DAG).
A C-DAG $\mathcal{G}_{c}$ is well-specified for $\mathcal{C}$ if it is induced by some well-specified $\mathcal{C}$-SCM-Prior.
The following theorem provides a necissary condition on whether a C-DAG can be well-specified and generalizes our insights from Lemma 4.2.
Theorem 4.8.
If $\mathcal{G}_{c}$ is well-specified for $\mathcal{C}$ and there is an unobserved confounder $U$ between $A$ and $Y$ in $\mathcal{G}_{c}$, then, under the regularity assumptions A.1 in Appendix A, either $A$ or $Y$ must have a noise variable as parent in $\mathcal{G}_{c}$.
Proof.
See Appendix A. ∎
The theorem yields a practical design rule: leave at least one exogenous-noise variable on the treatment–outcome path. Doing so lets the prior span both confounded and unconfounded worlds, a prerequisite for consistent posterior causal inference.
4.3 Constructing SCM-based priors
Figure 1: C-Dags compatible with the three example causal inference settings. Yellow variables are observed, blue variables are unobserved, red variables are clusters.
We now propose a practical algorithm to construct $\mathcal{C}$-SCM-priors. We assume that we have access to a pair $(\mathcal{G}_{c},\mathcal{I})$, where $\mathcal{G}_{c}$ is a well-specified C-DAG for $\mathcal{C}$ and $\mathcal{I}$ is a set of constraints on SCMs $\mathcal{S}$ compatible with $\mathcal{G}_{c}$ ensuring that $\mathcal{S}$ is also compatible with $\mathcal{C}$.
Example 1: back-door adjustment. The observable variables are $(X,A,Y)$ together with noise variables. A compatible C-DAG is in Fig. 1 (left). The constraint set is
$\mathcal{I}(\mathcal{S})\;=\;\bigl{\{}\mathbb{P}_{\mathrm{obs}}^{\mathcal{S}}(A=a\mid X=x)>0\bigr{\}},$ (19)
ensuring that all SCMs satisfy the positivity assumption.
Example 2: Front-door adjustment. Here, the obsesrved variables are $(X,A,M,Y)$ with noise variables and an unobserved confounder $U$ between $A$ and $Y$. A compatible C-DAG is in Fig. 1 (middle). The constraint set is
$\mathcal{I}(\mathcal{S})\;=\;\Bigl{\{}\mathbb{P}_{\mathrm{obs}}^{\mathcal{S}}(A=a\mid X=x)>0,\mathbb{P}_{\mathrm{obs}}^{\mathcal{S}}(M=m\mid X=x,A=a)>0\Bigr{\}},$ (20)
ensuring positivity for both treatments and mediators.
Example 3: Instrumental variables. The observed variables are $(X,Z,A,Y)$, augmented by noise variables and an unobserved confounder $U$ that is a joint parent of $A$ and $Y$ has no edge to the instrument $Z$; see the compatible C-DAG in Fig. 1 (right). The constraint set is
$\mathcal{I}(\mathcal{S})\;=\;\Bigl{\{}\mathbb{P}_{\mathrm{obs}}^{\mathcal{S}}(Z=z\mid X=x)>0,f_{Y}^{\mathcal{S}}(X,A,U)=f(X,A)+g(X,U)\Bigr{\}}$ (21)
Overall algorithm. Given $(\mathcal{G}_{c},\mathcal{I})$ we propose to construct a prior distribution $\Pi$ over SCMs as follows: First, we order the clusters $(C_{1},\dots,C_{k})$ according to their hierarchy in the DAG (i.e., $C_{1}$ has no parents). Then, we iterate over each cluster $C_{i}$ as follows: if $C_{i}$ only contains latent variables, fix their distribution to a standard normal distribution via
$U^{(i)}\;\sim\;\mathcal{N}\!\bigl{(}0,\mathbf{I}\bigr{)}.$ (22)
If $C_{i}$ is a cluster of observed and latent variables, we assign a clustered BNN prior to $C_{i}$ (see below). If $C_{i}$ only contains observed variables, we assign a observational BNN prior.
Clustered BNN prior. For clusters that contain both observed and latent variables, we leverage a BNN-based prior inspired by TabPFN [Hollmann.2023]. This prior allows us to effectively sample potentially high-dimensional clusters of variables for which the internal causal structure is irrelevant to infer the causal query of interest. The prior is defined via
$g_{\theta}^{(i)}:\;\mathrm{pa}(C_{i})\;\longrightarrow\;\mathbb{R}^{r},\qquad\theta\sim\Pi_{C_{i}}\quad\text{subject to }g_{\theta}^{(i)}\text{ satisfying }\mathcal{I}(\mathcal{S}_{\theta}).$ (23)
We then sample random nodes from $g_{\theta}^{(i)}$ that coincide with observed nodes in $C_{i}$, while the remaining nodes serve as latent noise within the cluster. This corresponds to applying the approach taking in TabPFN [Hollmann.2023] to clusters $C_{i}$ in the C-DAG in which the causal structure does not matter for estimating our causal query.
Observational BNN prior. If $C_{i}$ contains only observed nodes, we define another BNN via
$f_{\theta}^{(i)}:\;\mathrm{pa}(C_{i})\;\longrightarrow\;\mathbb{R}^{|C_{i}|},\qquad\theta\sim\Pi_{C_{i}}\quad\text{subject to }f_{\theta}^{(i)}\text{ satisfying }\mathcal{I}(\mathcal{S}_{\theta}),$ (24)
and set $C_{i}=f_{\theta}^{(i)}\!\bigl{(}\mathrm{pa}(C_{i})\bigr{)}$. The observed nodes within $C_{i}$ thus correspond to the output of the neural network and are not randomly subsampled neurons.
Example 1: back-door adjustment. Here, the data distribution $\mathbb{P}$ can be seperated as follows: $(X,U_{X})\sim\mathbb{P}_{X}$ with $U_{X}$ denoting noise variables withing the cluster $X$, $U_{A}\sim\mathbb{P}_{U_{A}}$, $U_{Y}\sim\mathbb{P}_{U_{Y}}$, $A=f_{A}(X,U_{A})$, and $Y=f_{A}(X,A,U_{Y})$. Our algorithm proceeds as follows: $\mathbb{P}_{U_{A}}$ and $\mathbb{P}_{U_{Y}}$ are noise variables and are set to standard normal distributions. The cluster $(X,U_{X})$ contains both noise and observed variables, meaning that $\mathbb{P}_{X}$ is sampled from an clustered BNN prior. Finally, $A$ and $Y$ are observed variables meaning that $f_{A}$ and $f_{Y}$ are sampled from observational BNN priors.
Implementation details. (i) For each covariate cluster $C_{i}$ containing latent nodes, we sample a random MLP-style graph over $\mathrm{pa}\left(C_{i}\right)$ by drawing biases and edge-weights from $\Pi_{C_{i}}$ and then pruning edges at random to ensure acyclicity. We evaluate this graph with tanh activations and noise (from normal, uniform, Laplace, or logistic distribution) to produce continuous features, then apply randomized thresholds to discretize or binarize a subset, yielding mixed-type covariates via our unstructured BNN prior. (ii) For treatment (and outcome) clusters $C_{j}$ of purely observed nodes, we instantiate a second BNN $f_{\theta}^{(j)}$ over $\mathrm{pa}\left(C_{j}\right)$ (with $\theta\sim\Pi_{C_{j}}$ and the same acyclicity constraint). We forward-propagate the covariates through $f_{\theta}^{(j)}$ with injected noise to compute a scalar propensity score, then threshold to assign a binary treatment. We forward-propagate both covariates and treatment to obtain potential outcomes. The resulting treatment (and outcome) are sampled from our structured BNN prior.
5 CausalFM: Training
5.1 Training algorithm
We look at the case where the causal query $Q(\mathbb{P}_{\mathrm{int}}(Y\mid X))$ is a function of the conditional interventional distribution $\mathbb{P}_{\mathrm{int}}(Y\mid X)$ for some contextual observed variables $X$. This includes, e.g., the CATE $\mathbb{E}[Y(1)-Y(0)\mid X]$ and CAPO $\mathbb{E}[Y(a)\mid X]$ from our running examples.
Our goal is to train a PFN $q_{\theta}(Y\mid x)$ to approximate the conditional PPID
$\Pi_{\mathrm{int}}(Y\mid\mathcal{D}_{n},X=x)=\int\mathbb{P}_{\mathrm{int}}^{\mathcal{S}}(Y\mid X=x)\,\Pi(\mathcal{S}\mid\mathcal{D}_{n})\,\mathop{}\!\mathrm{d}\mathcal{S}.$ (25)
Given an SCM prior $\Pi$ and a prior $\Pi_{N}$ over sample sizes, we propose the following modified PFN loss
$\mathcal{L}(\theta)=\mathbb{E}_{N\sim\Pi_{N}}\!\bigl{[}\mathbb{E}_{\mathcal{S}\sim\Pi}\bigl{[}\mathbb{E}_{(X,Y)\sim\mathbb{P}_{\mathrm{int}}^{\mathcal{S}}}\bigl{[}\mathbb{E}_{\mathcal{D}\sim\mathbb{P}_{\mathrm{obs}}^{\mathcal{S}}}\bigl{[}-\log q_{\theta}(Y\mid X,\mathcal{D}_{N})\bigr{]}\bigr{]}\bigr{]}\bigr{]}.$ (26)
Importantly, the dataset $\mathcal{D}$ is sampled from the observational distribution, while the pair $(X,Y)$ is sampled from the interventional distribution induced by a random SCM. This ensures that the PFN will aim to predict the interventional outcome $Y$ based on data following the observational distribution. A similar loss has been proposed by [Bynum.2025], which, however, is only based on the mean-squared error instead of the negative log-likelihood and thus does not allow an interpretation for approximating the PPID in a Bayesian setting. In particular, modeling the entire PPID allows us not only to provide point estimators of our causal query, but also account for uncertainty.
In practice, we sample a sample size $N_{j}\sim\Pi_{N}$, an SCM $\mathcal{S}_{j}\sim\Pi$ and an observational dataset $\mathcal{D}_{N_{j}}^{j}\sim\mathbb{P}_{\mathrm{obs}}^{\mathcal{S}_{j}}$ by sampling from the SCM. Then, we modify the SCM by performing the intervention of interest (e.g., $do(A=a)$) and sample test points $(x_{j},y_{j})\sim\mathbb{P}_{\mathrm{int}}^{\mathcal{S}_{j}}$ from the interventional SCM. The approximated PFN-loss is then
^$\hat{\mathcal{L}}(\theta)=\sum_{j}\bigl{[}-\log q_{\theta}(y_{j}\mid\mathcal{D}_{N_{j}}^{j},x_{j})\bigr{]}.$ (27)
Finally, once $q_{\theta}(Y\mid x)$ is trained, we can obtain an estimator for the causal query via $Q(q_{\theta}(Y\mid X)$, i.e., by applying the causal query on the approximated PPID by the PFN.
Example: back-door adjustment.
Here, sample a sample size $N_{j}\sim\Pi_{N}$, an SCM $\mathcal{S}_{j}\sim\Pi$ from our constructed prior distribution $\Pi$ and an observational dataset $\mathcal{D}_{N_{j}}^{j}\sim\mathbb{P}_{\mathrm{obs}}^{\mathcal{S}_{j}}$. Then, we perform two interventions $do(A=1)$ and $do(A=0)$ to obtain test points $(x_{j},y_{j}(1)-y_{j}(0))\sim\mathbb{P}_{\mathrm{int}}^{\mathcal{S}_{j}}$. The PFN loss becomes
^$\hat{\mathcal{L}}(\theta)=\sum_{j}\bigl{[}-\log q_{\theta}(y_{j}(1)-y_{j}(0)\mid\mathcal{D}_{N_{j}}^{j},x_{j})\bigr{]}.$ (28)
Model architecture and training details. Our model implementation of $q_{\theta}(Y\mid x)$ builds upon the TabPFN architecture [Hollmann.2023]. Each observation in the dataset is tokenized during the embedding process. We use separate encoders for covariates, treatment, and outcome. These embedded tokens are processed through a transformer where attention is applied between the observations. The model uses a task-specific decoder to predict CATE. The classification head consists of a two-layer MLP with GELU activation. The model is trained with cross-entropy loss with learning rate $1\mathrm{e}{-3}$, weight decay $1\mathrm{e}{-5}$), batch size $16$, and sequence length $1024$. Training proceeds for up to $100$ epochs with early stopping after $10$ epochs of no improvement on the validation loss. We generate $10k$ synthetic datasets for training. Training our CausalFM model on an NVIDIA A100 GPU takes approximately $10$ hours. The full implementation details are in Appendix B.
6 Experiments
6.1 Set up
Baselines. We consider a broad array of state-of-the-art methods for treatment effect estimation from the literature. Considering the rich methods in the literature, we use the following baselines: (1) S-learner [kunzel2019metalearners]: the S-learner is a model-agnostic learner that trains a single regression model by concatenating the covariate and the treatment as input; (2) T-learner [kunzel2019metalearners]: the T-learner is a model-agnostic learner that trains separate regression models for treated and control groups; (3) X-learner [kunzel2019metalearners]: builds upon the T-learner by first imputing individual treatment effects in each group and then fitting models to these pseudo-effects—often yielding more accurate estimates when treatment and control group sizes are imbalanced. (4) TARNet [shalit2017estimating]: using representation learning to extract features of covariates and train separate branches for treated and control groups with regularization.(5) DR-learner [kennedy2023towards]: generates pseudo-outcomes based on the doubly-robust AIPW estimator; (6) RA-learner [curth2021nonparametric]: uses a regression-adjusted pseudo-outcome in the second stage. Further implementation details are in Appendix B.1.
Evaluation metrics. We use the precision of estimating the heterogeneous effects (PEHE) criterion [curth2021nonparametric, hill2011bayesian] to evaluate model performance on the CATE estimation task.
Table 1: Results for benchmarking CATE evaluation performance across 10 datasets. Reported: mean $\pm$ standard deviation over five-fold train-test splits.
$\mathrm{D}_{1}$ $\mathrm{D}_{2}$ $\mathrm{D}_{3}$ $\mathrm{D}_{4}$ $\mathrm{D}_{5}$ $\mathrm{D}_{6}$ $\mathrm{D}_{7}$ $\mathrm{D}_{8}$ $\mathrm{D}_{9}$ $\mathrm{D}_{10}$
S-learner [kunzel2019metalearners] 0.265 $\pm$ 0.005 0.375 $\pm$ 0.006 0.438 $\pm$ 0.005 0.577 $\pm$ 0.005 0.562 $\pm$ 0.005 0.726 $\pm$ 0.004 0.537 $\pm$ 0.007 0.472 $\pm$ 0.005 0.652 $\pm$ 0.008 0.685 $\pm$ 0.004
T-learner [kunzel2019metalearners] 0.271 $\pm$ 0.008 0.385 $\pm$ 0.010 0.458 $\pm$ 0.008 0.602 $\pm$ 0.012 0.587 $\pm$ 0.006 0.728 $\pm$ 0.004 0.623 $\pm$ 0.008 0.535 $\pm$ 0.011 0.581 $\pm$ 0.006 0.626 $\pm$ 0.010
TARNet [shalit2017estimating] 0.347 $\pm$ 0.024 0.727 $\pm$ 0.023 0.536 $\pm$ 0.023 0.844 $\pm$ 0.026 1.203 $\pm$ 0.018 0.782 $\pm$ 0.020 1.379 $\pm$ 0.018 0.929 $\pm$ 0.021 0.959 $\pm$ 0.019 1.211 $\pm$ 0.022
DR-learner [kennedy2023towards] 0.351 $\pm$ 0.013 0.558 $\pm$ 0.016 0.599 $\pm$ 0.012 0.734 $\pm$ 0.019 0.703 $\pm$ 0.015 0.781$\pm$ 0.016 0.726 $\pm$ 0.016 0.612 $\pm$ 0.010 0.781 $\pm$ 0.018 0.799 $\pm$ 0.016
RA-learner [curth2021nonparametric] 0.315 $\pm$ 0.015 0.407 $\pm$ 0.014 0.459 $\pm$ 0.013 0.546 $\pm$ 0.016 0.620 $\pm$ 0.011 0.641 $\pm$ 0.018 0.608 $\pm$ 0.017 0.522 $\pm$ 0.013 0.602 $\pm$ 0.012 0.655 $\pm$ 0.017
X-learner [kunzel2019metalearners] 0.280 $\pm$ 0.128 0.393 $\pm$ 0.124 0.458 $\pm$ 0.126 0.562 $\pm$ 0.018 0.602 $\pm$ 0.124 0.642 $\pm$ 0.130 0.609 $\pm$ 0.130 0.506 $\pm$ 0.127 0.628 $\pm$ 0.126 0.646 $\pm$ 0.125
CausalFM (ours) 0.263 $\pm$ 0.002 0.333 $\pm$ 0.001 0.417 $\pm$ 0.001 0.541 $\pm$ 0.003 0.559 $\pm$ 0.001 0.641 $\pm$ 0.001 0.508 $\pm$ 0.001 0.426 $\pm$ 0.001 0.561 $\pm$ 0.001 0.620 $\pm$ 0.002
Lower = better (best in bold).
Table 2: Results showing the CATE evaluation performance on the Jobs.
PEHE
S-learner [kunzel2019metalearners] 0.598 $\pm$ 0.004
T-learner [kunzel2019metalearners] 0.618 $\pm$ 0.008
TARNet [shalit2017estimating] 1.305 $\pm$ 0.025
DR-learner [kennedy2023towards] 0.663 $\pm$ 0.016
RA-learner [curth2021nonparametric] 0.611 $\pm$ 0.014
X-learner [kunzel2019metalearners] 0.801 $\pm$ 0.129
CausalFM (ours) 0.581 $\pm$ 0.002
Lower = better (best in bold).
Synthetic dataset. Due to the fundamental problem of causal inference, the ground-truth CATE can not be observed in real-world data. We thus generate synthetic datasets to evaluate model performance following prior literature (e.g., [kunzel2019metalearners, shalit2017estimating, kennedy2023towards]). We benchmark our model on ten datasets with different data generation mechanisms. Further details are provided in the Appendix B.
Jobs dataset. The study by LaLonde [lalonde1986evaluating] is a widely used benchmark in the causal inference field. This dataset combines a randomized study based on the National Supported Work program with observational data [smith2005does]. The treatment is job training and the outcomes are income and employment status after training. The study includes 8 covariates, such as age and education, as well as previous earnings. We use the Jobs dataset constructed by [shalit2017estimating], where the goal is to predict unemployment.
6.2 Results
Table 2 shows the benchmarking results on the Jobs dataset. Table 1 shows the benchmarking results on 10 synthetic datasets. The experiments demonstrate that our CausalFM achieves competitive CATE estimation performance on these datasets, without the need to re-train the model.
7 Conclusion and Future Work
This paper introduces CausalFM, a framework for training foundation models for causal inference using prior-data fitted networks (PFNs). While we instantiate the approach for estimating CATEs, this represents only a first step. The next steps are instantiations to broader causal inference settings like instrumental variables and front-door adjustment.
We are currently expanding our work, and, in the next step, we will instantiate other versions of CausalFM to broader causal inference settings as described above. In addition, we plan to scale CausalFM to larger and more diverse datasets to demonstrate the generalizability and practical applicability of our methods. Thereby, our aim to establish robust foundation models for a wide range of causal inference settings.
8 Acknowledgments
This work has been supported by the German Federal Ministry of Education and Research (Grant: 01IS24082).
[…]
Appendix A Proofs
A.1 Linear IV
Proof of Lemma 4.1.
We prove that for any $\zeta\neq\zeta^{*}$, there exists an SCM $S=(\alpha,\beta,\gamma,\delta,\zeta,\eta,\theta,\kappa\neq 0)$ that induces the same observational distribution as $S^{*}$.
Step 1: Observational Distribution of $S^{*}$
The observational distribution is characterized by the covariance matrix $\Sigma^{*}$ of $(Z^{*},A^{*},Y^{*})$:
$\mathrm{Var}(Z^{*})$ $=(\alpha^{*})^{2}$ (29)
$\text{Cov}(Z^{*},A^{*})$ $=(\alpha^{*})^{2}\beta^{*}$ (30)
$\mathrm{Var}(A^{*})$ $=(\alpha^{*}\beta^{*})^{2}+(\delta^{*})^{2}+(\gamma^{*})^{2}$ (31)
$\text{Cov}(Z^{*},Y^{*})$ $=\zeta^{*}(\alpha^{*})^{2}\beta^{*}$ (32)
$\text{Cov}(A^{*},Y^{*})$ $=\zeta^{*}[(\alpha^{*}\beta^{*})^{2}+(\delta^{*})^{2}+(\gamma^{*})^{2}]+\eta^{*}\gamma^{*}$ (33)
$\mathrm{Var}(Y^{*})$ $=\zeta^{*2}[(\alpha^{*}\beta^{*})^{2}+(\delta^{*})^{2}+(\gamma^{*})^{2}]+2\zeta^{*}\eta^{*}\gamma^{*}+(\eta^{*})^{2}+(\theta^{*})^{2}$ (34)
Step 2: Construction of Alternative SCM $S$
The covariance matrix $\Sigma$ has elements:
$\mathrm{Var}(Z)$ $=\alpha^{2}+\kappa^{2}$ (35)
$\text{Cov}(Z,A)$ $=\alpha^{2}\beta+\kappa(\kappa\beta+\gamma)$ (36)
$\mathrm{Var}(A)$ $=(\alpha\beta)^{2}+(\kappa\beta+\gamma)^{2}+\delta^{2}$ (37)
$\text{Cov}(Z,Y)$ $=\zeta(\alpha^{2}\beta+\kappa(\kappa\beta+\gamma))+\eta\kappa$ (38)
$\text{Cov}(A,Y)$ $=\zeta[(\alpha\beta)^{2}+(\kappa\beta+\gamma)^{2}+\delta^{2}]+\eta(\kappa\beta+\gamma)$ (39)
$\mathrm{Var}(Y)$ $=\zeta^{2}[(\alpha\beta)^{2}+(\kappa\beta+\gamma)^{2}+\delta^{2}]+2\zeta\eta(\kappa\beta+\gamma)+\eta^{2}+\theta^{2}$ (40)
Step 3: Parameter Matching
To achieve $\Sigma=\Sigma^{*}$, we need:
$\alpha^{2}+\kappa^{2}$ $=(\alpha^{*})^{2}$ (41)
$\alpha^{2}\beta+\kappa(\kappa\beta+\gamma)$ $=(\alpha^{*})^{2}\beta^{*}$ (42)
$(\alpha\beta)^{2}+(\kappa\beta+\gamma)^{2}+\delta^{2}$ $=(\alpha^{*}\beta^{*})^{2}+(\delta^{*})^{2}+(\gamma^{*})^{2}$ (43)
$\zeta(\alpha^{2}\beta+\kappa(\kappa\beta+\gamma))+\eta\kappa$ $=\zeta^{*}(\alpha^{*})^{2}\beta^{*}$ (44)
$\zeta[(\alpha\beta)^{2}+(\kappa\beta+\gamma)^{2}+\delta^{2}]+\eta(\kappa\beta+\gamma)$ $=\zeta^{*}[(\alpha^{*}\beta^{*})^{2}+(\delta^{*})^{2}+(\gamma^{*})^{2}]+\eta^{*}\gamma^{*}$ (45)
$\zeta^{2}[(\alpha\beta)^{2}+(\kappa\beta+\gamma)^{2}+\delta^{2}]+2\zeta\eta(\kappa\beta+\gamma)+\eta^{2}+\theta^{2}$ $=(\mathrm{Var}(Y^{*}))$ (46)
Step 4: Solution Construction
Choose $\kappa\neq 0$ such that $|\kappa|<|\alpha^{*}|$. Set:
$\alpha$ $=\sqrt{(\alpha^{*})^{2}-\kappa^{2}}$ (47)
[…]
$\delta$ $=\delta^{*}$ (49)
[…]
Since $\alpha\beta=\alpha\beta^{*}=\frac{\alpha}{\alpha^{*}}\alpha^{*}\beta^{*}$, we have:
$(\alpha\beta)^{2}=\frac{\alpha^{2}}{(\alpha^{*})^{2}}(\alpha^{*}\beta^{*})^{2}=\frac{(\alpha^{*})^{2}-\kappa^{2}}{(\alpha^{*})^{2}}(\alpha^{*}\beta^{*})^{2}$
Therefore:
$\kappa\beta+\gamma=\pm\sqrt{(\gamma^{*})^{2}+\frac{\kappa^{2}}{(\alpha^{*})^{2}}(\alpha^{*}\beta^{*})^{2}}$
From equations (44) and (45), we can solve for $\eta$:
$\eta=\frac{\zeta^{*}(\alpha^{*})^{2}\beta^{*}-\zeta(\alpha^{2}\beta+\kappa(\kappa\beta+\gamma))}{\kappa}$
Finally, $\theta$ is determined from equation (46).
Step 5: Existence Verification
The system has 8 parameters $(\alpha,\beta,\gamma,\delta,\zeta,\eta,\theta,\kappa)$ and 6 constraints (the 6 unique entries of the covariance matrix). Since $\zeta\neq\zeta^{*}$ is fixed and $\kappa\neq 0$ is chosen, we have 6 remaining parameters for 6 constraints. The key observation is that the introduction of confounding ($\kappa\neq 0$) creates additional correlation structures that can compensate for the change in the causal effect $\zeta$, allowing the observational distribution to remain unchanged.
Proof of Lemma 4.2..
Let $\mathcal{S}=(\alpha,\beta,\gamma,\delta,\zeta,\eta,\theta)$ be any SCM from the linear IV class in Eq. (14). Then the following coefficients are identified via obserational data alone:
$\alpha=\sqrt{\mathrm{Var}(Z)},\beta=\mathbb{E}[Y\mid Z=1],\zeta=\frac{\zeta\beta}{\beta}=\frac{\mathbb{E}[\zeta(\beta+\delta\epsilon_{A})]-\mathbb{E}[\zeta(\delta\epsilon_{A})]}{\beta}=\frac{\mathbb{E}[Y\mid Z=1]-\mathbb{E}[Y\mid Z=0]}{\mathbb{E}[A|Z=1]-\mathbb{E}[A|Z=0]}$ (51)
as well as the combination of coefficients
$\delta^{2}+\gamma^{2}=\mathrm{Var}(A\mid Z),\eta^{2}+\theta^{2}=\mathrm{Var}(Y\mid A)$ (52)
and the back-door adjustment
$\mathbb{E}[Y\mid A=1]-E[Y\mid A=0]$ $=\zeta+\eta(\mathbb{E}[U\mid A=1]-\mathbb{E}[U\mid A=0])$ (53)
$=\zeta+\frac{\eta\gamma}{\gamma^{2}+\delta^{2}+(\beta\alpha)^{2}}.$ (54)
Note that the back-door adjustment is biased for $\zeta$ due to unobserved confounding.
Noiseless treatment case. Let now $\mathcal{S}^{\ast}=(\alpha^{\ast},\beta^{\ast},\gamma^{\ast},\delta^{\ast},\zeta^{\ast},\eta^{\ast},\theta^{\ast})$ denote an arbitrary fixed SCM from the linear IV class. We start by constructing $\mathcal{S}_{1}=(\alpha_{1},\beta_{1},\gamma_{1},\delta_{1}=0,\zeta_{1},\eta_{1},\theta_{1})$ such that $\mathbb{P}_{\mathrm{obs}}^{\mathcal{S}_{1}}=\mathbb{P}_{\mathrm{obs}}^{\mathcal{S}^{\ast}}$. Because of Eq. (51), we set
$\alpha_{1}=\alpha^{\ast},\,\beta_{1}=\beta^{\ast},\,\zeta_{1}=\zeta^{\ast}.$ (55)
Furthermore, setting $\delta_{1}=0$ implies due to Eq. (52) that
$\gamma_{1}^{2}={\delta^{\ast}}^{2}+{\gamma^{\ast}}^{2}.$ (56)
Due to Eq. (53) it must holds that $\frac{\eta_{1}\gamma_{1}}{\gamma_{1}^{2}+(\beta_{1}\alpha_{1})^{2}}=\frac{\eta^{\ast}\gamma^{\ast}}{{\delta^{\ast}}^{2}+{\gamma^{\ast}}^{2}+(\beta^{\ast}\alpha^{\ast})^{2}}$, which implies that
$\eta_{1}^{2}=\frac{{\eta^{\ast}}^{2}{\gamma^{\ast}}^{2}}{{\delta^{\ast}}^{2}+{\gamma^{\ast}}^{2}}.$ (57)
Finally, due to Eq. (52),
$\theta_{1}^{2}={\eta^{\ast}}^{2}+{\theta^{\ast}}^{2}-\frac{{\eta^{\ast}}^{2}{\gamma^{\ast}}^{2}}{{\delta^{\ast}}^{2}+{\gamma^{\ast}}^{2}},$ (58)
which means that every parameter of $\mathcal{S}_{1}$ has a unique solution in terms of parameters of $\mathcal{S}^{\ast}$ under the constraints of preserving the observational distribution.
Noiseless outcome case. We now construct $\mathcal{S}_{2}=(\alpha_{2},\beta_{2},\gamma_{2},\delta_{2},\zeta_{2},\eta_{2},\theta_{2}=0)$ such that $\mathbb{P}_{\mathrm{obs}}^{\mathcal{S}_{2}}=\mathbb{P}_{\mathrm{obs}}^{\mathcal{S}^{\ast}}$. Again, Eq. (51) implies that
$\alpha_{2}=\alpha^{\ast},\,\beta_{2}=\beta^{\ast},\,\zeta_{2}=\zeta^{\ast},$ (59)
and setting $\theta_{2}=0$ implies due to Eq. (52) that
$\eta_{2}^{2}={\eta^{\ast}}^{2}+{\theta^{\ast}}^{2}.$ (60)
Due to Eq. (53) it must holds that $\frac{\eta\gamma}{\gamma^{2}+\delta^{2}+(\beta\alpha)^{2}}=\frac{\eta\gamma}{{\delta^{\ast}}^{2}+{\gamma^{\ast}}^{2}+(\beta^{\ast}\alpha^{\ast})^{2}}$ which implies that
$\gamma_{2}^{2}=\frac{{\eta^{\ast}}^{2}{\gamma^{\ast}}^{2}}{{\eta^{\ast}}^{2}+{\theta^{\ast}}^{2}}.$ (61)
Finally, due to Eq. (52),
$\delta^{2}={\eta^{\ast}}^{2}+{\gamma^{\ast}}^{2}-\frac{{\eta^{\ast}}^{2}{\gamma^{\ast}}^{2}}{{\eta^{\ast}}^{2}+{\theta^{\ast}}^{2}},$ (62)
which means that every parameter of $\mathcal{S}_{2}$ has a unique solution in terms of parameters of $\mathcal{S}^{\ast}$ under the constraints of preserving the observational distribution. ∎
A.2 Theorem 4.8
We strart by imposing the following regularity assumption on the causal inference setting $\mathcal{C}$.
Assumption A.1 (Regularity assumptions).
  1. 1.
    Positivity and non-degeneracy: all admissible $\mathbb{P}_{\mathrm{obs}}\in\mathcal{P}_{\mathrm{obs}}$ satisfy $\mathrm{Var}_{\mathbb{P}_{\mathrm{obs}}}\!\bigl{(}A\mid pa(A)_{\mathrm{obs}}\bigr{)}>0$ and $\mathrm{Var}_{\mathbb{P}_{\mathrm{obs}}}\!\bigl{(}Y\mid pa(Y)_{\mathrm{obs}}\bigr{)}>0$.
  2. 2.
    Possibility of unconfoundedness: there exists an SCM $\mathcal{S}$ compatible with $\mathcal{C}$ in which $A$ and $Y$ are unconfounded.
  3. 3.
    Confounding changes the query: a confounded SCM $\mathcal{S}_{1}$ and an unconfounded $\mathcal{S}_{2}$ both compatible with $\mathcal{C}$ cannot coincide in their observational distributions, i.e., $\mathbb{P}_{\mathrm{obs}}^{S_{1}}\neq\mathbb{P}_{\mathrm{obs}}^{S_{2}}$.
Point 1) is a classical positivity assumption as common in the causal inference literature. Point 2) ensures that an unconfounded DGP is a possibility when generating the observational data. Finally, point 3) ensures that causal query $Q$ is affected by the unobserved confounder $U$ between $A$ and $Y$. If 3) would not hold, we find an unconfounded $\mathcal{S}_{1}$ and and unconfounded $\mathcal{S}_{2}$ with $\mathbb{P}_{\mathrm{obs}}^{\mathcal{S}_{1}}=\mathbb{P}_{\mathrm{obs}}^{\mathcal{S}_{2}}$, which also implies $Q(\mathbb{P}_{\mathrm{int}}^{\mathcal{S}_{1}})=\bar{Q}(\mathbb{P}_{\mathrm{obs}}^{\mathcal{S}_{1}})=\bar{Q}(\mathbb{P}_{\mathrm{obs}}^{\mathcal{S}_{2}})=Q(\mathbb{P}_{\mathrm{int}}^{\mathcal{S}_{2}})$ by identifiability of the causal query. Hence, this assumption is realistic for various causal queries, such as CAPO or CATE.
Proof.
Let $\Pi$ be a well-specified $\mathcal{C}$-SCM-prior with induced C-DAG $G_{c}$ that has an unobserved confounder between $A$ and $Y$. We have to show that in $G_{c}$ either $A$ or $Y$ has a noise parent.
By Assumption A.1 (2), there exists SCM $\mathcal{S^{\ast}}$ compatible with $\mathcal{C}$, in which $A$ and $Y$ are unconfounded. Let $\mathcal{D}_{n}\sim\mathbb{P}_{\mathrm{int}}^{\mathcal{S}^{\ast}}$ be sampled from the unconfounded observational distribution. Because $\Pi$ is well specified for $\mathcal{C}$, it holds that
$Q\!\Bigl{(}\int\mathbb{P}_{\mathrm{int}}^{\mathcal{S}}\,\Pi(\mathcal{S}\mid\mathcal{D}_{n})\,\mathop{}\!\mathrm{d}\mathcal{S}\Bigr{)}\;\longrightarrow\;Q(\mathbb{P}_{\mathrm{int}}^{\mathcal{S}^{\ast}}),\quad n\to\infty.$ (63)
However, this must imply that for all $n\in\mathbb{N}$ there exists an SCM $\mathcal{S}$ with posterior $\Pi(\mathcal{S}\mid\mathcal{D}_{n})>0$, and hence $\Pi(\mathcal{S})>0$ and $\Pi(\mathcal{D}_{n}\mid\mathcal{S})>0$. Thus in the limit for $n\to\infty$ this implies $\mathbb{P}_{\mathrm{obs}}^{\mathcal{S}}=\mathbb{P}_{\mathrm{obs}}^{\mathcal{S}^{\ast}}$. By assumption  A.1 (3) this implies that $\mathcal{S}$ is unconfounded. Hence, because of Assumption A.1 (1), either $A$ or $Y$ must have a noise parent in $\mathcal{S}$. Because $\Pi(\mathcal{S}>0)$, either $A$ or $Y$ must have also a noise parent in $G_{c}$.
Appendix B Implementation details
We implement our CausalFM using PyTorch. Our transformer-based model is designed for conditional average treatment effect (CATE) estimation. Our model implementation builds upon the TabPFN architecture [Hollmann.2023] https://github.com/PriorLabs/TabPFN/tree/main. We use separate encoders for each modality. Each observation in the dataset is tokenized during the embedding process. The embedded tokens are then processed through a basic transformer where attention is applied between the observations. In the final layer, tokens corresponding to the inference observations are passed through a classification head. We frame CATE estimation as a three-class classification task by discretizing the individual treatment effect into three categories: negative effect, no effect, and positive effect. The model uses a task-specific decoder to predict CATE. During training, we use cross-entropy loss computed on model output and target classes. When training from scratch, the model is optimized using the Adam optimizer with a learning rate of $31e-$ 3 and weight decay of $51e-$ 5 .. The training batch size is set to $16$ and the sequence length is set to $1024$. The total training epoch is $100$. Early stopping is triggered after $10$ epochs of no improvement in validation loss. We generate $10$ k datasets. Empirically, training our causalFM model takes about $10$ hours on an NVIDIA A100 GPU.
B.1 Implementation details of baselines
We follow the implementation from https://github.com/AliciaCurth/CATENets/tree/main for most of the CATE estimators, including S-learner [kunzel2019metalearners], T-learner [kunzel2019metalearners], TARNet [shalit2017estimating], X-leaner [kunzel2019metalearners], DR-learner [kennedy2023towards], RA-learner [curth2021nonparametric]. For each dataset and method, we evaluated 5 repetitions, each with a different random seed. All methods used the same train-test split.